# Hallucination Detection in LLMs Using Spectral Features of Attention Maps

### v1

@@ D.2 Best found hyperparameters
We present the hyperparameter values corresponding to the results in Table 1 and Table 4. Table 5 shows the optimal hyperparameter $k$ for selecting the top-$k$ eigenvalues from either the attention maps in $\operatorname{AttnEigvals}$ or the Laplacian matrix in $\operatorname{LapEigvals}$. While fewer eigenvalues were sufficient for optimal performance in some cases, the best results were generally achieved with the highest tested value, $k=100$.
@@
LLM Temp Feature AllLayers PerLayer top-$k$ eigenvalues
CoQA HaluevalQA NQOpen SQuADv2 TriviaQA TruthfulQA
Llama3.1-8B 0.1 $\operatorname{AttnEigvals}$ 50 100 25 100 100 10
Llama3.1-8B 0.1 $\operatorname{AttnEigvals}$ 100 100 100 100 50 100
Llama3.1-8B 0.1 $\operatorname{LapEigvals}$ 50 100 10 100 100 100
Llama3.1-8B 0.1 $\operatorname{LapEigvals}$ 10 100 100 100 100 100
Llama3.1-8B 1.0 $\operatorname{AttnEigvals}$ 100 100 100 100 100 100
Llama3.1-8B 1.0 $\operatorname{AttnEigvals}$ 100 100 100 100 100 100
Llama3.1-8B 1.0 $\operatorname{LapEigvals}$ 100 100 100 100 100 100
Llama3.1-8B 1.0 $\operatorname{LapEigvals}$ 100 25 100 100 100 100
Llama3.2-3B 0.1 $\operatorname{AttnEigvals}$ 100 100 100 100 100 10
Llama3.2-3B 0.1 $\operatorname{AttnEigvals}$ 100 25 100 100 100 100
Llama3.2-3B 0.1 $\operatorname{LapEigvals}$ 100 100 100 100 50 5
Llama3.2-3B 0.1 $\operatorname{LapEigvals}$ 25 100 100 100 100 100
Llama3.2-3B 1.0 $\operatorname{AttnEigvals}$ 100 100 100 100 100 50
Llama3.2-3B 1.0 $\operatorname{AttnEigvals}$ 100 100 100 100 100 100
Llama3.2-3B 1.0 $\operatorname{LapEigvals}$ 100 100 10 100 100 25
Llama3.2-3B 1.0 $\operatorname{LapEigvals}$ 25 100 100 100 100 100
Phi3.5 0.1 $\operatorname{AttnEigvals}$ 100 100 100 100 100 100
Phi3.5 0.1 $\operatorname{AttnEigvals}$ 100 10 10 25 100 50
Phi3.5 0.1 $\operatorname{LapEigvals}$ 100 100 100 100 100 100
Phi3.5 0.1 $\operatorname{LapEigvals}$ 10 50 100 100 100 100
Phi3.5 1.0 $\operatorname{AttnEigvals}$ 100 100 100 100 100 100
Phi3.5 1.0 $\operatorname{AttnEigvals}$ 100 100 10 100 100 50
Phi3.5 1.0 $\operatorname{LapEigvals}$ 100 100 100 100 100 50
Phi3.5 1.0 $\operatorname{LapEigvals}$ 10 100 100 100 100 100
@@
LLM $temp$ Feature Layer index
CoQA HaluevalQA NQOpen SQuADv2 TriviaQA TruthfulQA
Llama3.1-8B 0.1 $\operatorname{AttentionScore}$ 13 10 0 0 0 28
Llama3.1-8B 0.1 $\operatorname{AttnLogDet}$ 7 13 16 11 29 21
Llama3.1-8B 0.1 $\operatorname{AttnEigvals}$ 22 31 26 31 31 7
Llama3.1-8B 0.1 $\operatorname{LapEigvals}$ 15 14 20 29 31 20
Llama3.1-8B 1.0 $\operatorname{AttentionScore}$ 29 10 0 0 0 23
Llama3.1-8B 1.0 $\operatorname{AttnLogDet}$ 17 11 13 29 29 30
Llama3.1-8B 1.0 $\operatorname{AttnEigvals}$ 22 31 31 31 31 31
Llama3.1-8B 1.0 $\operatorname{LapEigvals}$ 15 14 31 29 29 29
Llama3.2-3B 0.1 $\operatorname{AttentionScore}$ 15 12 12 12 21 14
Llama3.2-3B 0.1 $\operatorname{AttnLogDet}$ 12 13 24 10 25 14
Llama3.2-3B 0.1 $\operatorname{AttnEigvals}$ 27 14 14 25 27 17
Llama3.2-3B 0.1 $\operatorname{LapEigvals}$ 11 8 12 25 12 14
Llama3.2-3B 1.0 $\operatorname{AttentionScore}$ 24 12 0 24 21 14
Llama3.2-3B 1.0 $\operatorname{AttnLogDet}$ 12 26 23 25 25 12
Llama3.2-3B 1.0 $\operatorname{AttnEigvals}$ 11 27 25 25 27 10
Llama3.2-3B 1.0 $\operatorname{LapEigvals}$ 11 18 12 25 25 11
Phi3.5 0.1 $\operatorname{AttentionScore}$ 7 15 0 0 0 19
Phi3.5 0.1 $\operatorname{AttnLogDet}$ 20 18 16 17 13 23
Phi3.5 0.1 $\operatorname{AttnEigvals}$ 18 19 15 19 18 28
Phi3.5 0.1 $\operatorname{LapEigvals}$ 18 28 28 19 31 28
Phi3.5 1.0 $\operatorname{AttentionScore}$ 19 0 1 0 0 19
Phi3.5 1.0 $\operatorname{AttnLogDet}$ 12 29 14 19 13 14
Phi3.5 1.0 $\operatorname{AttnEigvals}$ 18 30 17 31 31 31
Phi3.5 1.0 $\operatorname{LapEigvals}$ 18 28 15 19 31 31
Appendix E Extended results of generalisation study
@@
We present the complete results of the generalisation ablation discussed in Section 6.4 of the main paper. Table 7 reports the absolute Test AUROC values for each method and test dataset. Except for TruthfulQA, $\operatorname{LapEigvals}$ achieves the highest performance across all configurations. Notably, some methods perform close to random, whereas $\operatorname{LapEigvals}$ consistently outperforms this baseline. Regarding relative performance drop (Figure 10), $\operatorname{LapEigvals}$ remains competitive, exhibiting the lowest drop in nearly half of the scenarios. These results indicate that our method is robust but warrants further investigation across more datasets, particularly with a deeper analysis of TruthfulQA.
@@
Table 7: Full results of the generalisation study. By gray color we denote results obtained on test split from the same QA dataset as training split, otherwise results are from test split of different QA dataset. We highlight the best performance in bold.
@@
CoQA HaluevalQA NQOpen SQuADv2 TriviaQA TruthfulQA
$\operatorname{AttnLogDet}$ CoQA 0.758 0.687 0.644 0.646 0.640 0.587
$\operatorname{AttnEigvals}$ CoQA 0.782 0.726 0.696 0.659 0.702 0.560
$\operatorname{LapEigvals}$ CoQA 0.830 0.790 0.748 0.743 0.786 0.629
$\operatorname{AttnLogDet}$ HaluevalQA 0.580 0.823 0.750 0.727 0.787 0.668
$\operatorname{AttnEigvals}$ HaluevalQA 0.579 0.819 0.792 0.743 0.803 0.688
$\operatorname{LapEigvals}$ HaluevalQA 0.685 0.873 0.796 0.778 0.848 0.595
$\operatorname{AttnLogDet}$ NQOpen 0.552 0.720 0.794 0.717 0.766 0.597
$\operatorname{AttnEigvals}$ NQOpen 0.546 0.725 0.790 0.714 0.770 0.618
$\operatorname{LapEigvals}$ NQOpen 0.656 0.792 0.827 0.748 0.843 0.564
$\operatorname{AttnLogDet}$ SQuADv2 0.553 0.716 0.774 0.746 0.757 0.658
$\operatorname{AttnEigvals}$ SQuADv2 0.576 0.730 0.737 0.768 0.760 0.711
$\operatorname{LapEigvals}$ SQuADv2 0.673 0.801 0.806 0.791 0.841 0.625
$\operatorname{AttnLogDet}$ TriviaQA 0.565 0.761 0.793 0.736 0.838 0.572
$\operatorname{AttnEigvals}$ TriviaQA 0.577 0.770 0.786 0.742 0.843 0.616
$\operatorname{LapEigvals}$ TriviaQA 0.702 0.813 0.818 0.773 0.889 0.522
$\operatorname{AttnLogDet}$ TruthfulQA 0.550 0.597 0.603 0.604 0.662 0.811
$\operatorname{AttnEigvals}$ TruthfulQA 0.538 0.600 0.595 0.646 0.685 0.833
$\operatorname{LapEigvals}$ TruthfulQA 0.590 0.552 0.529 0.569 0.631 0.829

### v2

@@ G.3 Best found hyperparameters
We present the hyperparameter values corresponding to the results in Table 1 and Table 6. Table 8 shows the optimal hyperparameter $k$ for selecting the top-$k$ eigenvalues from either the attention maps in $\operatorname{AttnEigvals}$ or the Laplacian matrix in $\operatorname{LapEigvals}$. While fewer eigenvalues were sufficient for optimal performance in some cases, the best results were generally achieved with the highest tested value, k=100k{=}100.
@@
LLM Temp Feature all-layers per-layer top-$k$ eigenvalues
CoQA GSM8K HaluevalQA NQOpen SQuADv2 TriviaQA TruthfulQA
Llama3.1-8B 0.1 $\operatorname{AttnEigvals}$ 50 100 100 25 100 100 10
Llama3.1-8B 0.1 $\operatorname{AttnEigvals}$ 100 100 100 100 100 50 100
Llama3.1-8B 0.1 $\operatorname{LapEigvals}$ 50 50 100 10 100 100 100
Llama3.1-8B 0.1 $\operatorname{LapEigvals}$ 10 100 100 100 100 100 100
Llama3.1-8B 1.0 $\operatorname{AttnEigvals}$ 100 100 100 100 100 100 100
Llama3.1-8B 1.0 $\operatorname{AttnEigvals}$ 100 100 100 100 100 100 100
Llama3.1-8B 1.0 $\operatorname{LapEigvals}$ 100 50 100 100 100 100 100
Llama3.1-8B 1.0 $\operatorname{LapEigvals}$ 100 100 25 100 100 100 100
Llama3.2-3B 0.1 $\operatorname{AttnEigvals}$ 100 100 100 100 100 100 10
Llama3.2-3B 0.1 $\operatorname{AttnEigvals}$ 100 100 25 100 100 100 100
Llama3.2-3B 0.1 $\operatorname{LapEigvals}$ 100 25 100 100 100 50 5
Llama3.2-3B 0.1 $\operatorname{LapEigvals}$ 25 100 100 100 100 100 100
Llama3.2-3B 1.0 $\operatorname{AttnEigvals}$ 100 100 100 100 100 100 50
Llama3.2-3B 1.0 $\operatorname{AttnEigvals}$ 100 50 100 100 100 100 100
Llama3.2-3B 1.0 $\operatorname{LapEigvals}$ 100 50 100 10 100 100 25
Llama3.2-3B 1.0 $\operatorname{LapEigvals}$ 25 100 100 100 100 100 100
Phi3.5 0.1 $\operatorname{AttnEigvals}$ 100 100 100 100 100 100 100
Phi3.5 0.1 $\operatorname{AttnEigvals}$ 100 25 10 10 25 100 50
Phi3.5 0.1 $\operatorname{LapEigvals}$ 100 10 100 100 100 100 100
Phi3.5 0.1 $\operatorname{LapEigvals}$ 10 100 50 100 100 100 100
Phi3.5 1.0 $\operatorname{AttnEigvals}$ 100 100 100 100 100 100 100
Phi3.5 1.0 $\operatorname{AttnEigvals}$ 100 100 100 10 100 100 50
Phi3.5 1.0 $\operatorname{LapEigvals}$ 100 25 100 100 100 100 50
Phi3.5 1.0 $\operatorname{LapEigvals}$ 10 25 100 100 100 100 100
Mistral-Nemo 0.1 $\operatorname{AttnEigvals}$ 100 50 100 100 100 100 100
Mistral-Nemo 0.1 $\operatorname{AttnEigvals}$ 100 50 100 100 100 100 100
Mistral-Nemo 0.1 $\operatorname{LapEigvals}$ 100 25 100 100 100 100 10
Mistral-Nemo 0.1 $\operatorname{LapEigvals}$ 10 100 25 100 50 100 100
Mistral-Nemo 1.0 $\operatorname{AttnEigvals}$ 100 100 100 100 100 100 100
Mistral-Nemo 1.0 $\operatorname{AttnEigvals}$ 100 100 100 100 100 50 100
Mistral-Nemo 1.0 $\operatorname{LapEigvals}$ 100 100 100 50 100 100 100
Mistral-Nemo 1.0 $\operatorname{LapEigvals}$ 10 100 50 100 100 100 100
Mistral-Small-24B 0.1 $\operatorname{AttnEigvals}$ 100 100 100 10 100 50 25
Mistral-Small-24B 0.1 $\operatorname{AttnEigvals}$ 100 100 100 100 100 100 25
Mistral-Small-24B 0.1 $\operatorname{LapEigvals}$ 100 50 100 50 100 100 10
Mistral-Small-24B 0.1 $\operatorname{LapEigvals}$ 25 100 100 100 100 10 100
Mistral-Small-24B 1.0 $\operatorname{AttnEigvals}$ 100 100 100 100 100 100 100
Mistral-Small-24B 1.0 $\operatorname{AttnEigvals}$ 100 100 100 100 100 100 100
Mistral-Small-24B 1.0 $\operatorname{LapEigvals}$ 100 100 100 100 50 100 50
Mistral-Small-24B 1.0 $\operatorname{LapEigvals}$ 10 100 50 10 10 100 50
@@
LLM temptemp Feature Layer index
CoQA GSM8K HaluevalQA NQOpen SQuADv2 TriviaQA TruthfulQA
Llama3.1-8B 0.1 $\operatorname{AttentionScore}$ 13 28 10 0 0 0 28
Llama3.1-8B 0.1 $\operatorname{AttnLogDet}$ 7 31 13 16 11 29 21
Llama3.1-8B 0.1 $\operatorname{AttnEigvals}$ 22 31 31 26 31 31 7
Llama3.1-8B 0.1 $\operatorname{LapEigvals}$ 15 25 14 20 29 31 20
Llama3.1-8B 1.0 $\operatorname{AttentionScore}$ 29 3 10 0 0 0 23
Llama3.1-8B 1.0 $\operatorname{AttnLogDet}$ 17 16 11 13 29 29 30
Llama3.1-8B 1.0 $\operatorname{AttnEigvals}$ 22 28 31 31 31 31 31
Llama3.1-8B 1.0 $\operatorname{LapEigvals}$ 15 11 14 31 29 29 29
Llama3.2-3B 0.1 $\operatorname{AttentionScore}$ 15 17 12 12 12 21 14
Llama3.2-3B 0.1 $\operatorname{AttnLogDet}$ 12 18 13 24 10 25 14
Llama3.2-3B 0.1 $\operatorname{AttnEigvals}$ 27 14 14 14 25 27 17
Llama3.2-3B 0.1 $\operatorname{LapEigvals}$ 11 24 8 12 25 12 14
Llama3.2-3B 1.0 $\operatorname{AttentionScore}$ 24 25 12 0 24 21 14
Llama3.2-3B 1.0 $\operatorname{AttnLogDet}$ 12 18 26 23 25 25 12
Llama3.2-3B 1.0 $\operatorname{AttnEigvals}$ 11 14 27 25 25 27 10
Llama3.2-3B 1.0 $\operatorname{LapEigvals}$ 11 10 18 12 25 25 11
Phi3.5 0.1 $\operatorname{AttentionScore}$ 7 1 15 0 0 0 19
Phi3.5 0.1 $\operatorname{AttnLogDet}$ 20 19 18 16 17 13 23
Phi3.5 0.1 $\operatorname{AttnEigvals}$ 18 18 19 15 19 18 28
Phi3.5 0.1 $\operatorname{LapEigvals}$ 18 23 28 28 19 31 28
Phi3.5 1.0 $\operatorname{AttentionScore}$ 19 1 0 1 0 0 19
Phi3.5 1.0 $\operatorname{AttnLogDet}$ 12 19 29 14 19 13 14
Phi3.5 1.0 $\operatorname{AttnEigvals}$ 18 1 30 17 31 31 31
Phi3.5 1.0 $\operatorname{LapEigvals}$ 18 16 28 15 19 31 31
Mistral-Nemo 0.1 $\operatorname{AttentionScore}$ 2 27 18 35 0 30 35
Mistral-Nemo 0.1 $\operatorname{AttnLogDet}$ 37 20 17 15 38 38 33
Mistral-Nemo 0.1 $\operatorname{AttnEigvals}$ 38 37 38 18 18 15 31
Mistral-Nemo 0.1 $\operatorname{LapEigvals}$ 16 38 37 37 18 37 8
Mistral-Nemo 1.0 $\operatorname{AttentionScore}$ 10 2 16 28 14 30 21
Mistral-Nemo 1.0 $\operatorname{AttnLogDet}$ 18 17 20 18 18 15 18
Mistral-Nemo 1.0 $\operatorname{AttnEigvals}$ 38 30 39 39 18 15 18
Mistral-Nemo 1.0 $\operatorname{LapEigvals}$ 16 39 37 37 18 37 18
Mistral-Small-24B 0.1 $\operatorname{AttentionScore}$ 14 1 39 33 35 0 30
Mistral-Small-24B 0.1 $\operatorname{AttnLogDet}$ 16 29 38 18 16 38 11
Mistral-Small-24B 0.1 $\operatorname{AttnEigvals}$ 36 27 36 19 16 38 20
Mistral-Small-24B 0.1 $\operatorname{LapEigvals}$ 21 3 35 24 36 35 34
Mistral-Small-24B 1.0 $\operatorname{AttentionScore}$ 15 1 1 0 1 0 30
Mistral-Small-24B 1.0 $\operatorname{AttnLogDet}$ 14 24 27 17 24 38 34
Mistral-Small-24B 1.0 $\operatorname{AttnEigvals}$ 36 39 27 21 24 36 23
Mistral-Small-24B 1.0 $\operatorname{LapEigvals}$ 21 39 36 16 21 35 34
G.4 Comparison with hidden-states-based baselines
We take an approach considered in the previous works Azaria and Mitchell (2023); Orgad et al. (2025) and aligned to our evaluation protocol. Specifically, we trained a logistic regression classifier on PCA-projected hidden states to predict whether the model is hallucinating or not. To this end, we select the last token of the answer. While we also tested the last token of the prompt, we observed significantly lower performance, which aligns with results presented by (Orgad et al., 2025). We considered hidden states either from all layers or a single layer corresponding to the selected token. In the all-layer scenario, we use the concatenation of hidden states of all layers, and in the per-layer scenario, we use the hidden states of each layer separately and select the best-performing layer.
In Table 10, we show the obtained results. The all-layer version is consistently worse than our $\operatorname{LapEigvals}$, which further confirms the strength of the proposed method. Our work is one of the first to detect hallucinations solely using attention maps, providing an important insight into the behavior of LLMs, and it motivates further theoretical research on information flow patterns inside these models.
Table 10: Results of the probe trained on the hidden state features from the last generated token.
LLM Temp Features per-layer all-layers Test AUROC ($\uparrow$)
CoQA GSM8K HaluevalQA NQOpen SQuADv2 TriviaQA TruthfulQA
Llama3.1-8B 0.1 $\operatorname{HiddenStates}$ 0.835 0.799 0.840 0.766 0.736 0.820 0.834
Llama3.1-8B 0.1 $\operatorname{HiddenStates}$ 0.821 0.765 0.825 0.728 0.723 0.791 0.785
Llama3.1-8B 0.1 $\operatorname{LapEigvals}$ 0.757 0.844 0.793 0.711 0.733 0.780 0.764
Llama3.1-8B 0.1 $\operatorname{LapEigvals}$ 0.836 0.887 0.867 0.793 0.782 0.872 0.822
Llama3.1-8B 1.0 $\operatorname{HiddenStates}$ 0.836 0.816 0.850 0.786 0.754 0.850 0.823
Llama3.1-8B 1.0 $\operatorname{HiddenStates}$ 0.835 0.759 0.847 0.757 0.749 0.838 0.808
Llama3.1-8B 1.0 $\operatorname{LapEigvals}$ 0.743 0.833 0.789 0.725 0.724 0.794 0.764
Llama3.1-8B 1.0 $\operatorname{LapEigvals}$ 0.830 0.872 0.874 0.827 0.791 0.889 0.829
Llama3.2-3B 0.1 $\operatorname{HiddenStates}$ 0.800 0.826 0.808 0.732 0.750 0.782 0.760
Llama3.2-3B 0.1 $\operatorname{HiddenStates}$ 0.790 0.802 0.784 0.709 0.721 0.760 0.770
Llama3.2-3B 0.1 $\operatorname{LapEigvals}$ 0.676 0.835 0.774 0.730 0.727 0.712 0.690
Llama3.2-3B 0.1 $\operatorname{LapEigvals}$ 0.801 0.852 0.844 0.771 0.778 0.821 0.743
Llama3.2-3B 1.0 $\operatorname{HiddenStates}$ 0.778 0.727 0.758 0.679 0.719 0.773 0.716
Llama3.2-3B 1.0 $\operatorname{HiddenStates}$ 0.773 0.652 0.753 0.657 0.681 0.761 0.618
Llama3.2-3B 1.0 $\operatorname{LapEigvals}$ 0.715 0.815 0.765 0.696 0.696 0.738 0.767
Llama3.2-3B 1.0 $\operatorname{LapEigvals}$ 0.812 0.870 0.857 0.798 0.751 0.836 0.787
Phi3.5 0.1 $\operatorname{HiddenStates}$ 0.841 0.773 0.845 0.813 0.781 0.886 0.737
Phi3.5 0.1 $\operatorname{HiddenStates}$ 0.833 0.696 0.840 0.806 0.774 0.878 0.689
Phi3.5 0.1 $\operatorname{LapEigvals}$ 0.716 0.753 0.757 0.761 0.732 0.768 0.741
Phi3.5 0.1 $\operatorname{LapEigvals}$ 0.810 0.785 0.819 0.815 0.791 0.858 0.717
Phi3.5 1.0 $\operatorname{HiddenStates}$ 0.872 0.784 0.850 0.821 0.806 0.891 0.822
Phi3.5 1.0 $\operatorname{HiddenStates}$ 0.853 0.686 0.844 0.804 0.790 0.887 0.752
Phi3.5 1.0 $\operatorname{LapEigvals}$ 0.723 0.816 0.769 0.755 0.732 0.792 0.732
Phi3.5 1.0 $\operatorname{LapEigvals}$ 0.821 0.885 0.836 0.826 0.795 0.872 0.777
Mistral-Nemo 0.1 $\operatorname{HiddenStates}$ 0.818 0.757 0.814 0.734 0.731 0.821 0.792
Mistral-Nemo 0.1 $\operatorname{HiddenStates}$ 0.805 0.741 0.784 0.722 0.730 0.793 0.699
Mistral-Nemo 0.1 $\operatorname{LapEigvals}$ 0.759 0.751 0.760 0.697 0.696 0.769 0.710
Mistral-Nemo 0.1 $\operatorname{LapEigvals}$ 0.823 0.805 0.821 0.755 0.767 0.858 0.737
Mistral-Nemo 1.0 $\operatorname{HiddenStates}$ 0.793 0.832 0.777 0.738 0.719 0.783 0.722
Mistral-Nemo 1.0 $\operatorname{HiddenStates}$ 0.771 0.834 0.771 0.706 0.685 0.779 0.644
Mistral-Nemo 1.0 $\operatorname{LapEigvals}$ 0.738 0.808 0.763 0.708 0.723 0.785 0.818
Mistral-Nemo 1.0 $\operatorname{LapEigvals}$ 0.835 0.890 0.833 0.795 0.812 0.865 0.828
Mistral-Small-24B 0.1 $\operatorname{HiddenStates}$ 0.838 0.872 0.744 0.680 0.700 0.749 0.735
Mistral-Small-24B 0.1 $\operatorname{HiddenStates}$ 0.815 0.812 0.703 0.632 0.629 0.726 0.589
Mistral-Small-24B 0.1 $\operatorname{LapEigvals}$ 0.800 0.850 0.719 0.674 0.784 0.757 0.827
Mistral-Small-24B 0.1 $\operatorname{LapEigvals}$ 0.852 0.881 0.808 0.722 0.821 0.831 0.757
Mistral-Small-24B 1.0 $\operatorname{HiddenStates}$ 0.801 0.879 0.720 0.665 0.603 0.684 0.581
Mistral-Small-24B 1.0 $\operatorname{HiddenStates}$ 0.770 0.760 0.703 0.617 0.575 0.659 0.485
Mistral-Small-24B 1.0 $\operatorname{LapEigvals}$ 0.805 0.897 0.790 0.712 0.781 0.779 0.725
Mistral-Small-24B 1.0 $\operatorname{LapEigvals}$ 0.861 0.925 0.882 0.791 0.820 0.876 0.748
Appendix H Extended results of ablations
In the following section, we extend the ablation results presented in Section 6.1 and Section 6.2. Figure 10 compares the top $k$ eigenvalues across all five LLMs. In Figure 11 we present a layer-wise performance comparison for each model.
Figure 10: Probe performance across different top-$k$ eigenvalues: $k\in\{5,10,25,50,100\}$ for TriviaQA dataset with temp=1.0temp{=}1.0 and five considered LLMs.
Figure 11: Analysis of model performance across different layers for and 5 considered LLMs and TriviaQA dataset with temp=1.0temp{=}1.0 and k=100k{=}100 top eigenvalues (results for models operating on all layers provided for reference).
Appendix I Extended results of generalization study
@@
We present the complete results of the generalization ablation discussed in Section 6.4 of the main paper. Table 11 reports the absolute Test AUROC values for each method and test dataset. Except for TruthfulQA, $\operatorname{LapEigvals}$ achieves the highest performance across all configurations. Notably, some methods perform close to random, whereas $\operatorname{LapEigvals}$ consistently outperforms this baseline. Regarding relative performance drop (Figure 12), $\operatorname{LapEigvals}$ remains competitive, exhibiting the lowest drop in nearly half of the scenarios. These results indicate that our method is robust but warrants further investigation across more datasets, particularly with a deeper analysis of TruthfulQA.
@@
Table 11: Full results of the generalization study. By gray color we denote results obtained on test split from the same QA dataset as training split, otherwise results are from test split of different QA dataset. We highlight the best performance in bold.
@@
CoQA GSM8K HaluevalQA NQOpen SQuADv2 TriviaQA TruthfulQA
$\operatorname{AttnLogDet}$ CoQA 0.758 0.518 0.687 0.644 0.646 0.640 0.587
$\operatorname{AttnEigvals}$ CoQA 0.782 0.426 0.726 0.696 0.659 0.702 0.560
$\operatorname{LapEigvals}$ CoQA 0.830 0.555 0.790 0.748 0.743 0.786 0.629
$\operatorname{AttnLogDet}$ GSM8K 0.515 0.828 0.513 0.502 0.555 0.503 0.586
$\operatorname{AttnEigvals}$ GSM8K 0.510 0.838 0.563 0.545 0.549 0.579 0.557
$\operatorname{LapEigvals}$ GSM8K 0.568 0.872 0.648 0.596 0.611 0.610 0.538
$\operatorname{AttnLogDet}$ HaluevalQA 0.580 0.500 0.823 0.750 0.727 0.787 0.668
$\operatorname{AttnEigvals}$ HaluevalQA 0.579 0.569 0.819 0.792 0.743 0.803 0.688
$\operatorname{LapEigvals}$ HaluevalQA 0.685 0.448 0.873 0.796 0.778 0.848 0.595
$\operatorname{AttnLogDet}$ NQOpen 0.552 0.594 0.720 0.794 0.717 0.766 0.597
$\operatorname{AttnEigvals}$ NQOpen 0.546 0.633 0.725 0.790 0.714 0.770 0.618
$\operatorname{LapEigvals}$ NQOpen 0.656 0.676 0.792 0.827 0.748 0.843 0.564
$\operatorname{AttnLogDet}$ SQuADv2 0.553 0.695 0.716 0.774 0.746 0.757 0.658
$\operatorname{AttnEigvals}$ SQuADv2 0.576 0.723 0.730 0.737 0.768 0.760 0.711
$\operatorname{LapEigvals}$ SQuADv2 0.673 0.754 0.801 0.806 0.791 0.841 0.625
$\operatorname{AttnLogDet}$ TriviaQA 0.565 0.618 0.761 0.793 0.736 0.838 0.572
$\operatorname{AttnEigvals}$ TriviaQA 0.577 0.667 0.770 0.786 0.742 0.843 0.616
$\operatorname{LapEigvals}$ TriviaQA 0.702 0.612 0.813 0.818 0.773 0.889 0.522
$\operatorname{AttnLogDet}$ TruthfulQA 0.550 0.706 0.597 0.603 0.604 0.662 0.811
$\operatorname{AttnEigvals}$ TruthfulQA 0.538 0.579 0.600 0.595 0.646 0.685 0.833
$\operatorname{LapEigvals}$ TruthfulQA 0.590 0.722 0.552 0.529 0.569 0.631 0.829